OpenStudy (anonymous):

3costheta-sin(theta/2)-1=0

OpenStudy (cwrw238):

use the identity cos theta = 1 - 2 sin^2(theta/2)

OpenStudy (cwrw238):

let t = sin(theta/2) so 3(1 - t^2) - t - 1 = 0 3 - 3t^2 - t -1 = 0 3t^2 + t - 2 = 0 solve this for t then equate the solutions to sin(theta/2)

OpenStudy (anonymous):

youre a rock star thanks :)